how do i write y-3=1/2(x-4)in slope intercept form
Answer by Alan3354(69443)

how do i write y-3=1/2(x-4)in slope intercept form
--------------
Solve for y
y = 1/2(x-4) + 3
y = (1/2)x - 2 + 3
y = (1/2)x + 1